OpenStudy (griffindore):

A candle burns down at the rate of 0.5 inches per hour. The original height of the candle was 6 inches. Part A: Write a list of 6 ordered pairs to show the height of the candle in inches (y) as a function of time in hours (x) from the first hour after it started burning. For example, the point (0, 6) would represent a height of 6 inches after 0 hours. Explain how you obtained the ordered pairs. Part B: Is this relation a function? Justify your answer using the list of ordered pairs you created in Part A. Part C: If the rate at which the candle burnt was 0.3 inches per

OpenStudy (nick5006):

Do you need help on all of this?

OpenStudy (griffindore):

i mean... probably

OpenStudy (griffindore):

probably not

OpenStudy (nick5006):

Okei, with the information we know we could create an equation to show how long the candle will take to melt. With that we can plug in numbers to create ordered pairs

OpenStudy (griffindore):

ok

OpenStudy (nick5006):

So f(h)=6-(0.5h) will be the equation to show how long the candle will take to melt. "6" being the initial amount, 0.5 will be the rate, and "h" will be the amount of time in hours.

OpenStudy (griffindore):

ok

OpenStudy (nick5006):

H is also the X value, so all you have to do is plug in numbers from 0-6 to get the ordered pairs

OpenStudy (griffindore):

ok but wouldnt you have two different values? instead of two h's

OpenStudy (nick5006):

Okei for example f(0)=6-(0.5(0)) is f(0)=6. The "f(h)" is basically your "x" and the answer is your "y". So that would be your first ordered pair (0,6).

OpenStudy (nick5006):

All you have to do is plug in the rest of your numbers for more ordered pairs.

OpenStudy (griffindore):

then 1,6 and so on?

OpenStudy (nick5006):

Yup

OpenStudy (nick5006):

Actually wait no sorry

OpenStudy (griffindore):

oh ok...

OpenStudy (nick5006):

You have to actually plug it in and do the equation. f(h) is what you are inputing and the answer is the output

OpenStudy (nick5006):

f(1)=6-(.5(1)), what do you get?

OpenStudy (griffindore):

wait whats the f for?

OpenStudy (griffindore):

it would be5.5

OpenStudy (griffindore):

so its not a function

OpenStudy (nick5006):

It's a function, if you haven't gotten that far in math class I can explain it to you, but you need it in the problem. You basically read "f(h)", as "function of h". This "h" can be anything else like x or t.

OpenStudy (griffindore):

oh i have, im actually in algerbra 1 but this just confused me... long story

OpenStudy (nick5006):

I mean do you get what I'm explaining?

OpenStudy (griffindore):

yah

OpenStudy (nick5006):

This is a strong concept in math, which you'll be using later in advance stuff. So if you want to get a better grasp I would recomend watching a video online. There's plenty of stuff that could explain it better

OpenStudy (griffindore):

ok thx

OpenStudy (griffindore):

oh one more thing... the last part sas, would it continue being a function...

OpenStudy (nick5006):

Do you need anymore help on, like how to prove if it's a function

OpenStudy (nick5006):

Okei so there's many ways to prove if it's a function

OpenStudy (griffindore):

ik but its not...

OpenStudy (nick5006):

No it's a function

OpenStudy (griffindore):

ugghhh this is confusoing

OpenStudy (griffindore):

nvm

OpenStudy (nick5006):

Anything is a function, unless it has repeating "x" values. It can have repeating "y" values but not x. So you can graph the equation, and by using the "vertical line test" you can tell if it is or not. You just draw vertical lines on the graph and if you notice if it's touching two points of your equation it's not a function

OpenStudy (nick5006):

Any further explanation, or is this good?